# Behavior Foundation Model: Towards Next-Generation Whole-Body Control System of Humanoid Robots

### v1

@@ 3.1.3 Forward-backward Representation Learning
Equipped with the successor measure, the action-value function $Q^{\pi}_{r}(\bm{s},\bm{a})$ of $\pi$ for any reward function $r:\mathcal{S}\rightarrow\mathbb{R}$ satisfies
$Q_{r}^{\pi}(\bm{s},\bm{a})$ $\coloneqq\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r\left(\bm{s}_{t+1}\right)|\bm{s},\bm{a},\pi\right]$ (4)
@@
Figure 7: Motivo [41] learns broad behavior coverage and demonstrates outstanding zero-shot adaptation capability to a diverse downstream tasks, including complex motion imitation, pose reaching, and composite reward optimization. Moreover, Motivo achieves real-time motor control while ensuring motion naturalness.
@@
The FB framework provides a general and flexible approach for training BFMs by learning successor measure representations and applying pre-trained policies to new tasks. However, it suffers from several limitations: (i) when the latent dimension $d$ is finite, it relies on a low-rank dynamics assumption, leading to limited inductive bias for policy selection; (ii) poor coverage in the training dataset causes offline learning to fail in reliably optimizing policies, often collapsing to a few suboptimal behaviors with weak performance on downstream tasks. These limitations greatly hinder the application of the FB framework for humanoid robots. To address these limitations, [41] proposes FB with conditional policy regularization (FB-CPR) and introduces Motivo, the first BFM for humanoid WBC in the true sense that solves diverse tasks in a zero-shot manner, including motion tracking, goal reaching, and reward optimization.
@@
where $\mathcal{D}_{\rm online}$ is the associated replay buffer of unsupervised transitions, $p_{\pi}(\bm{s},\bm{z})$ is the joint distribution of $(\bm{s},\bm{z})$ induced by FB, and $p_{\mathcal{E}}$ is the joint distribution of the dataset. However, it is intractable to optimize the divergence term directly via a RL procedure. To tackle the problem, FB-CPR interprets the divergence as an expected return under the polices and defines a divergence-based reward $r_{\rm div}$ as
@@
It is natural to find that FB-CPR is not a rigorous unsupervised method, as it leverages unlabeled demo data to assist motion prior learning. By aligning unsupervised RL with human-like behavioral priors from unlabeled data, FB-CPR enhances policy diversity and dataset coverage, enabling the agent to learn a rich latent space of behaviors (e.g., walking, jumping, handstands) and achieve robust zero-shot performance across diverse tasks. Experimental results demonstrate that Motivo achieves 83% success rate in motion tracking tasks, and performs 61% of the top-line performance in reward optimization tasks, surpassing DIFFUSER in computational efficiency by requiring only 12 seconds per 300-step episode. Additionally, it outperforms ASE and CALM in motion diversity, achieving a score of 4.70 ($\pm$0.66), reflecting its ability to capture a broader range of behaviors.
@@ 3.2 Adaptation
Equipped with the derived BFMs through the aforementioned pre-training frameworks, we further introduce recent advancements on the adaptation techniques of BFMs. These work can be broadly categorized into two types: fine-tuning and towards hierarchical control.
@@ 3.2.1 Fine-tuning
Fine-tuning of BFMs seeks to bridge the gap between general-purpose motion priors and task-specific requirements. While pre-trained BFMs capture broad motion distributions, they often lack precision for specialized tasks or novel environments. For example, [42] introduces fast adaptation techniques like residual latent adaptation (ReLA) and lookahead latent adaptation (LoLA), which allow BFMs to achieve zero-shot adaptation by using minimal online interactions after pre-training. These methods improve task performance by up to 40%, demonstrating the effectiveness of fast adaptation strategies that enable efficient task switching without requiring extensive retraining. In contrast, [43] introduces ”Task Tokens”, which enhance goal-conditioned BFMs by generating task-specific tokens through a task encoder, enabling the BFM to perform complex tasks like motion tracking and goal-reaching with high success rates. Their method shows up to 99.75% success in tasks such as the long jump. Additionally, [44] addresses the limitation of FB representations by introducing belief-FB, which uses a transformer encoder to infer environmental dynamics, improving adaptation to unseen changes. Their approach achieves up to 2$\times$ improvement in performance under dynamic variations, demonstrating enhanced zero-shot capabilities. Future work could investigate additional post-training techniques, such as test-time scaling [89] or RL from human feedback [90], to further improve the adaptability and efficiency of BFMs and ensure alignment with human preferences in real-world applications.
@@ 3.2.2 Towards Hierarchical Control
While fine-tuning techniques enhance the performance of BFMs for specific tasks through minimal modifications during test-time, several pioneering works have attempted to establish a hierarchical control architecture based on BFM, which decouples high-level planning from low-level motion execution to achieve more scalable and flexible control [91, 92]. For example, UniHSI [93] introduces a unified framework for human-scene interaction by using language commands to guide a chain-of-contacts, which represents the sequence of human-object contact pairs. The system translates language inputs into structured task plans, which are then executed by a unified controller based on the AMP architecture. This framework achieves semantic alignment between language commands and physical motions, supporting diverse interactions with single or multiple objects. TokenHSI [94] further extends this line by proposing a transformer-based unified policy that tokenizes human proprioception and task states. By separating shared motor knowledge (proprioception token) from task-specific parameters (task tokens), it enables seamless multi-skill unification and flexible adaptation to novel tasks, such as skill composition (e.g., carrying while sitting), object or terrain shape variation, and long-horizon task completion.
In contrast, CLoSD [95] introduces a text-driven RL controller that combines motion diffusion models with physics-based simulations for robust multi-task human character control. By utilizing a real-time diffusion planner and a motion tracking controller in a closed-loop feedback system, CLoSD can handle complex tasks like goal-reaching, striking, and human-object interactions, all controlled through text prompts and target locations. Similarly, UniPhys [96] introduces a diffusion-based behavior cloning framework unifying planning and control with diffusion forcing to handle prediction errors, enabling flexible control via text, velocity, and goal guidance for applications like dynamic obstacle avoidance and long-horizon planning. These applications demonstrate that BFMs serve as a pivotal bridge between high-level semantic instructions and low-level physical execution, leveraging pre-trained behavioral priors to enable zero-shot adaptation, multi-task generalization, and physics-aware motion synthesis. By integrating various control paradigms (e.g., transformer-based tokenization and closed-loop diffusion planning), these highlight BFMs’ potential to democratize humanoid control across complex, real-world scenarios, from interactive robotics to dynamic environment adaptation.
Figure 8: An overview of the application, limitations, research opportunities, and potential risks of BFMs.
@@ 4 Applications and Limitations
BFMs are foreseen to significantly enhance humanoid robotics by providing a universal pre-trained controller capable of generalizing across diverse tasks. In this section, we explore the potential applications of BFMs in diverse industries macroscopically, such as healthcare robotics and gaming. Furthermore, we identify the key limitations of current BFMs, including the Sim2Real gap, data bottleneck, and embodiment generalization. Our analysis is inspired by the current development and applications of other foundation models, such as LLMs [97] and large vision models [98].
@@ 4.2.2 Data Bottleneck
The data bottleneck poses a fundamental constraint in developing BFMs for humanoid robots. While datasets listed in Table II have been successfully employed to train the current BFMs, their scale remains significantly smaller than the datasets used to train LLMs or large vision models. This scarcity is exacerbated when retargeting motions to specific robotic platforms [138], where subtle morphological differences can incur severe policy performance loss. Real-world robot data is even more constrained due to the hardware limitations and safety concerns. The challenge compounds when considering multimodal data requirements.
TABLE II: Available humanoid motion datasets for training BFMs.
@@
Current BFMs predominantly rely on proprioceptive inputs, integrating exteroceptive sensing for real-world deployment introduces new bottlenecks. Pioneering work like OmniH2O [147] explores multimodal data collection through teleoperation and pairing head-mounted RGBD cameras with whole-body proprioception and motor commands. While this enables tasks like vision-based autonomy and imitation learning, its 40-minute dataset (OmniH2O-6) remains limited to lab environments and lacks diverse contact dynamics. Currently, there is no existing dataset that provides large-scale and temporally aligned recordings of proprioception, vision, and contact dynamics across diverse environments. Therefore, larger-scale, high-quality, and more curated datasets are urgently needed to enhance the effectiveness of current BFMs and to aid in the development of future models.
@@ 5.1.2 High-level ML System
Recent advances in foundation models like LLMs and VLAs have demonstrated their potential as controllers for coordinating specialized AI systems [151, 152, 153, 29]. For example, HuggingGPT [151] leverages LLMs to orchestrate task planning across diverse models, while Eureka [29] employs LLMs to automate reward function design in RL. These projects present a natural pathway for integrating BFMs into high-level ML systems, where they could serve as universal low-level controllers for humanoid robots. By combining the reasoning capabilities of LLMs (for task planning and adaptation) with the motor priors encoded in BFMs (for real-time execution), such systems could achieve unprecedented flexibility in handling complex and multi-step physical tasks, while minimizing task-specific engineering. The ultimate vision is to create a unified cognitive-physical architecture based on LLMs and BFMs, mirroring the seamless integration of human cognition and physical control.
@@ 5.1.3 Scaling Law
The concept of scaling laws describes how the performance of neural language models improves with increases in model size, data, and compute resources. While these laws are well-established in domains like language and vision, their applicability to BFMs remains an open problem [154, 155, 156, 157, 158]. Preliminary evidence suggests that scaling BFMs through larger architectures, diverse training datasets, and expanded computational resources can enhance their generalization and zero-shot adaptation capabilities. For instance, FB-CPR [41] demonstrated improved performance in humanoid control tasks with more parameters (from 25M to 288M), achieving more robust zero-shot motion tracking and reward optimization. On the other hand, data scaling appears more critical for BFMs, as it directly determines the quality and physical plausibility of the learned behavior priors, the foundational requirements that model scaling alone cannot address. However, the effect of data scaling remains underexplored, with open questions as discussed in Section 4.2.2. Furthermore, unlike LLMs, BFM scaling must strike an appropriate balance between the behavior coverage (diverse motor skills) and control efficiency (precision and real-time stability). Future work should rigorously quantify these scaling dynamics to unlock BFMs’ full potential as general-purpose controllers.
@@ 5.1.4 Post-training
Post-training techniques have emerged as critical tools for the continued success and refinement of foundation models, especially for LLMs [159], which refine models to improve reasoning, address limitations, and better align with user intents and ethical considerations. Among these methods, fine-tuning [160, 161, 162, 163, 164, 165], integration of RL [90, 166, 167, 168, 169, 170, 171, 172], and test-time scaling [173, 174, 174, 175, 89] have been the most prominent strategies for optimizing LLMs’ performance. Integrating these post-training strategies presents unique research opportunities for BFMs. For instance, leveraging RL techniques like RLHF and RLAIF for BFMs can be crucial for refining the alignment between the agent’s behavior and real-world human expectations, especially in human-centric task environments. This opens avenues for developing more robust models that can adapt dynamically to user feedback. Additionally, test-time scaling for BFMs can optimize the computational efficiency during deployment, especially for real-time robot control or decision-making systems. Research could focus on improving the scalability of BFMs while ensuring that model outputs remain accurate and contextually appropriate across varying operational conditions.
@@ 5.1.5 Multi-agent System
A multi-agent system (MAS) consists of multiple interacting agents that can be either cooperative, competitive, or a mix of both, aiming to address complex tasks that require collaboration, decision-making, and behavior coordination among agents [176, 177, 178, 179, 180, 181]. BFMs can fundamentally accelerate the construction of MAS consists of humanoid robots, eliminating the need to effortlessly teach each robot basic survival skills like balance and locomotion before they can collaborate. Instead, researchers can focus directly on higher-level coordination challenges like role allocation and team strategy. However, current BFMs trained on single-robot data lack specialized interaction capabilities needed for optimal collaboration. This presents a promising research direction: developing next-generation BFMs trained explicitly on multi-robot interaction scenarios. Such models could better handle physical coordination challenges like object handovers, formation maintenance, and collision avoidance while preserving their generalizability.
@@ 5.1.6 Evaluation Mechanism
While foundation models like LLMs benefit from well-established benchmarks (e.g., GPQA [182] for broad knowledge recall, MATH [183] for mathematical problem solving, or MUSR [184] for multi-step reasoning), there is no specific and comprehensive evaluation mechanism for evaluating BFM’s capability and guiding the evolution direction. A robust assessment of BFMs must consider multiple interdependent factors, including task generalization across unseen scenarios, adaptability to new skills with minimal data, robustness against physical perturbations, and alignment with human safety and interpretability standards. For example, Motivo [41] combines quantitative metrics like task success rates with qualitative human evaluations of motion naturalness, yet critical gaps remain in assessing compositional skill combinations, hardware-specific constraints, and long-term behavioral stability. Future benchmarks should focus on progressive difficulty levels and cross-domain transfer tests to effectively assess BFMs’ potential as general-purpose physical controllers in both simulated and real-world scenarios. This approach will ultimately guide the field towards developing more capable and reliable humanoid systems.

### v2

@@ 3.1.3 Forward-backward Representation Learning
Equipped with the successor measure, the action-value function $Q^{\pi}_{r}(\bm{s},\bm{a})$ of $\pi$ for any reward function $r$ satisfies
$Q_{r}^{\pi}(\bm{s},\bm{a})\coloneqq$ $\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r\left(\bm{s}_{t+1}\right)|\bm{s},\bm{a},\pi\right]$ (5)
@@
Figure 7: Motivo [46] learns broad behavior priors and demonstrates outstanding zero-shot adaptation capability to diverse downstream tasks, including complex motion imitation, pose reaching, and composite reward optimization. Moreover, Motivo achieves real-time motor control while ensuring motion naturalness.
The FB framework provides a general and flexible approach for training BFMs by learning successor measure representations and applying pre-trained policies to new tasks. However, it suffers from several limitations: (i) when the latent dimension $d$ is finite, it relies on the assumption of a low-rank dynamic, leading to a limited inductive bias for policy selection; (ii) poor coverage in the training dataset causes offline learning to fail in reliably optimizing policies, often collapsing to a few suboptimal behaviors with weak performance on downstream tasks. These limitations greatly hinder the application of the FB framework for humanoid robots. To address these limitations, [46] proposes FB with conditional policy regularization (FB-CPR) and introduces Motivo, as illustrated in Fig. 7, which solves diverse humanoid WBC tasks in a zero-shot manner, including motion tracking, goal reaching, and reward optimization.
@@
where $\mathcal{D}_{\rm online}$ is the associated replay buffer of unsupervised transitions, $p_{\pi}(\bm{s},\bm{z})$ is the joint distribution of $(\bm{s},\bm{z})$ induced by FB, and $p_{\mathcal{E}}$ is the joint distribution of the dataset. However, it is intractable to optimize the divergence term directly via a RL procedure. To tackle the problem, FB-CPR interprets the divergence as an expected return under the policies and defines a divergence-based reward $r_{\rm div}$ as
@@
It is natural to find that FB-CPR is not a rigorous unsupervised method, as it leverages unlabeled demo data to assist motion prior learning. By aligning unsupervised RL with human-like behavioral priors from unlabeled data, FB-CPR enhances policy diversity and dataset coverage, enabling the agent to learn a rich latent space of behaviors (e.g., walking, jumping, handstands) and achieve robust zero-shot performance across diverse tasks. Experimental results demonstrate that Motivo achieves 83% success rate in motion tracking tasks, and performs 61% of the top-line performance in reward optimization tasks, surpassing DIFFUSER [101] in computational efficiency by requiring only 12 seconds per 300-step episode. Additionally, it outperforms ASE and CALM in motion diversity, achieving a score of 4.70 ($\pm$0.66), reflecting its ability to capture a broader range of behaviors.
@@ 3.2 Adaptation
Equipped with the derived BFMs through the aforementioned pre-training frameworks, we further introduce recent advancements in the adaptation techniques of BFMs. These work can be broadly categorized into two types: fine-tuning and towards hierarchical control.
@@ 3.2.1 Fine-tuning
Fine-tuning of BFMs seeks to bridge the gap between general-purpose motion priors and task-specific requirements. While pre-trained BFMs capture broad motion distributions, they often lack precision for specialized tasks or novel environments. For example, [47] introduces fast adaptation techniques like residual latent adaptation (ReLA) and lookahead latent adaptation (LoLA), which allow BFMs to achieve zero-shot adaptation by using minimal online interactions after pre-training. These methods improve task performance by up to 40%, demonstrating the effectiveness of fast adaptation strategies that enable efficient task switching without requiring extensive retraining. In contrast, [48] introduces TaskTokens, which enhance goal-conditioned BFMs by generating task-specific tokens through a task encoder, enabling the BFM to perform complex tasks like motion tracking and goal-reaching with high success rates. Their method shows up to 99.75% success in tasks such as the long jump. Additionally, [49] addresses the limitation of FB representations by introducing belief-FB, which uses a transformer encoder to infer environmental dynamics, improving adaptation to unseen changes. Their approach achieves up to 2$\times$ improvement in performance under dynamic variations, demonstrating enhanced zero-shot capabilities. Future work could investigate additional post-training techniques, such as test-time scaling [102] or RL from human feedback [103], to further improve the adaptability and efficiency of BFMs and ensure alignment with human preferences in real-world applications.
@@ 3.2.2 Towards Hierarchical Control
While fine-tuning techniques enhance the performance of BFMs for specific tasks through minimal modifications during test time, several pioneering works have attempted to establish a hierarchical control architecture based on BFM, which decouples high-level planning from low-level motion execution to achieve more scalable and flexible control [87, 88]. For example, UniHSI [84] introduces a unified framework for human-scene interaction by using language commands to guide a chain-of-contacts, which represents the sequence of human-object contact pairs. The system translates language inputs into structured task plans, which are then executed by a unified controller based on the AMP architecture. This framework achieves semantic alignment between language commands and physical motions, supporting diverse interactions with single or multiple objects. TokenHSI [83] further extends this line by proposing a transformer-based unified policy that tokenizes human proprioception and task states. By separating shared motor knowledge (proprioception token) from task-specific parameters (task tokens), it enables seamless multi-skill unification and flexible adaptation to novel tasks, such as skill composition (e.g., carrying while sitting), object or terrain shape variation, and long-horizon task completion.
In contrast, CLoSD [86] introduces a text-driven RL controller that combines motion diffusion models with physics-based simulations for robust multi-task human character control. By utilizing a real-time diffusion planner and a motion tracking controller in a closed-loop feedback system, CLoSD can handle complex tasks like goal-reaching, striking, and human-object interactions, all controlled through text prompts and target locations. Similarly, UniPhys [85] introduces a diffusion-based behavior cloning framework unifying planning and control with diffusion forcing to handle prediction errors, enabling flexible control via text, velocity, and goal guidance for applications like dynamic obstacle avoidance and long-horizon planning. These applications demonstrate that BFMs serve as a pivotal bridge between high-level semantic instructions and low-level physical execution, leveraging pre-trained behavioral priors to enable zero-shot adaptation, multi-task generalization, and physics-aware motion synthesis. By integrating various control paradigms (e.g., transformer-based tokenization, and closed-loop diffusion planning), these highlight BFMs’ potential to democratize humanoid control across complex, real-world scenarios, from interactive robotics to dynamic environment adaptation.
@@ 4 Applications and Limitations
BFMs are foreseen to significantly enhance humanoid robotics by providing a universal pre-trained controller capable of generalizing across diverse tasks. In this section, we explore the potential applications of BFMs in diverse industries macroscopically, such as healthcare robotics and gaming. Furthermore, we identify the key limitations of current BFMs, including the Sim2Real gap, data bottleneck, and embodiment generalization. Our analysis is inspired by the current development and applications of other foundation models, such as LLMs [104] and large vision models [105, 106].
Figure 8: An overview of the practical applications and current limitations of BFMs.
@@ 4.2.2 Data Bottleneck
The data bottleneck poses a fundamental constraint in developing BFMs for humanoid robots. While datasets listed in Table III have been successfully employed to train the current BFMs, their scale remains significantly smaller than the datasets used to train LLMs or large vision models. This scarcity is exacerbated when retargeting motions to specific robotic platforms [146], where subtle morphological differences can incur severe policy performance loss. Real-world robot data is even more constrained due to hardware limitations and safety concerns. The challenge compounds when considering multimodal data requirements.
TABLE III: Available humanoid motion datasets for training BFMs.
@@
Current BFMs predominantly rely on proprioceptive inputs, and integrating exteroceptive sensing for real-world deployment introduces new bottlenecks. Pioneering work like OmniH2O [155] explores multimodal data collection through teleoperation and pairing head-mounted RGBD cameras with whole-body proprioception and motor commands. While this enables tasks like vision-based autonomy and imitation learning, its 40-minute dataset (OmniH2O-6) remains limited to lab environments and lacks diverse contact dynamics. Currently, there is no existing dataset that provides large-scale and temporally aligned recordings of proprioception, vision, and contact dynamics across diverse environments. Therefore, larger-scale, high-quality, and more curated datasets are urgently needed to enhance the effectiveness of current BFMs and to aid in the development of future models.
@@ 5 Opportunities and Risks
Figure 9: An overview of the future research opportunities and potential risks of BFMs.
@@ 5.1.2 High-level ML System
Recent advances in foundation models like LLMs and VLAs have demonstrated their potential as controllers for coordinating specialized AI systems [159, 160, 161, 34, 162, 163, 164]. For example, HuggingGPT [159] leverages LLMs to orchestrate task planning across diverse models, while Eureka [34] employs LLMs to automate reward function design in RL. These projects present a natural pathway for integrating BFMs into high-level ML systems, where they could serve as universal low-level controllers for humanoid robots. By combining the reasoning capabilities of LLMs (for task planning and adaptation) with the motor priors encoded in BFMs (for real-time execution), such systems could achieve unprecedented flexibility in handling complex and multi-step physical tasks, while minimizing task-specific engineering. The ultimate vision is to create a unified cognitive-physical architecture based on LLMs and BFMs, mirroring the seamless integration of human cognition and physical control.
@@ 5.1.3 Scaling Law
The concept of scaling laws describes how the performance of neural language models improves with increases in model size, data, and compute resources. While these laws are well-established in domains like language and vision, their applicability to BFMs remains an open problem [165, 166, 167, 168, 169]. Preliminary evidence suggests that scaling BFMs through larger architectures, diverse training datasets, and expanded computational resources can enhance their generalization and zero-shot adaptation capabilities. For instance, FB-CPR [46] demonstrated improved performance in humanoid control tasks with more parameters (from 25M to 288M), achieving more robust zero-shot motion tracking and reward optimization. On the other hand, data scaling appears more critical for BFMs, as it directly determines the quality and physical plausibility of the learned behavior priors, the foundational requirements that model scaling alone cannot address. However, the effect of data scaling remains underexplored, with open questions as discussed in Section 4.2.2. Furthermore, unlike LLMs, BFM scaling must strike an appropriate balance between behavior coverage (diverse motor skills) and control efficiency (precision and real-time stability). Future work should rigorously quantify these scaling dynamics to unlock BFMs’ full potential as general-purpose controllers.
@@ 5.1.4 Post-training
Post-training techniques have emerged as critical tools for the continued success and refinement of foundation models, especially for LLMs [170], which refine models to improve reasoning, address limitations, and better align with user intents and ethical considerations. Among these methods, fine-tuning [171, 172, 173, 174, 175, 176, 177], integration of RL [103, 178, 179, 180, 181, 182, 183, 184], and test-time scaling [185, 186, 186, 187, 102] have been the most prominent strategies for optimizing LLMs’ performance. Integrating these post-training strategies presents unique research opportunities for BFMs. For instance, leveraging RL techniques like RLHF and RLAIF for BFMs can be crucial for refining the alignment between the agent’s behavior and real-world human expectations, especially in human-centric task environments. This opens avenues for developing more robust models that can adapt dynamically to user feedback. Additionally, test-time scaling for BFMs can optimize the computational efficiency during deployment, especially for real-time robot control or decision-making systems. Research could focus on improving the scalability of BFMs while ensuring that model outputs remain accurate and contextually appropriate across varying operational conditions.
@@ 5.1.5 Multi-agent System
A multi-agent system (MAS) consists of multiple interacting agents that can be either cooperative, competitive, or a mix of both, aiming to address complex tasks that require collaboration, decision-making, and behavior coordination among agents [188, 189, 190, 191, 192, 193]. BFMs can fundamentally accelerate the construction of MAS consists of humanoid robots, eliminating the need to effortlessly teach each robot basic survival skills like balance and locomotion before they can collaborate. Instead, researchers can focus directly on higher-level coordination challenges like role allocation and team strategy. However, current BFMs trained on single-robot data lack the specialized interaction capabilities needed for optimal collaboration. This presents a promising research direction: developing next-generation BFMs trained explicitly on multi-robot interaction scenarios. Such models could better handle physical coordination challenges like object handovers, formation maintenance, and collision avoidance while preserving their generalizability.
@@ 5.1.6 Evaluation Mechanism
While foundation models like LLMs benefit from well-established benchmarks (e.g., GPQA [194] for broad knowledge recall, MATH [195] for mathematical problem solving, or MUSR [196] for multi-step reasoning), there is no specific and comprehensive evaluation mechanism for evaluating BFM’s capability and guiding the evolution direction [197, 198]. A robust assessment of BFMs must consider multiple interdependent factors, including task generalization across unseen scenarios, adaptability to new skills with minimal data, robustness against physical perturbations, and alignment with human safety and interpretability standards. For example, Motivo [46] combines quantitative metrics like task success rates with qualitative human evaluations of motion naturalness, yet critical gaps remain in assessing compositional skill combinations, hardware-specific constraints, and long-term behavioral stability. Future benchmarks should focus on progressive difficulty levels and cross-domain transfer tests to effectively assess BFMs’ potential as general-purpose physical controllers in both simulated and real-world scenarios. This approach will ultimately guide the field toward developing more capable and reliable humanoid systems.